OpenStudy (anonymous):

Does anyone know how to do a problem by completing the square? I'm so confused.

OpenStudy (mertsj):

We all do.

OpenStudy (anonymous):

yes, we do

OpenStudy (anonymous):

but best to work with a specific example

OpenStudy (mertsj):

Even the Great Satellite stands ready to help you.

OpenStudy (anonymous):

example: 4x^2 + 8x + 1 Using a^2 + 2ab + b^2 = (a + b)^2, we have (2x + 1)^2 There's other formulas.

OpenStudy (anonymous):

like a^2 - 2ab + b^2 = (a - b)^2

OpenStudy (anonymous):

example:3x²-6x-9=0

OpenStudy (anonymous):

step number one is to divide by three and get \[x^2-2x-3=0\] then add 3 to both sides and get \[x^2-2x=3\] and now we are ready to "complete the square" although we can also solve this one by factoring

OpenStudy (anonymous):

half of -2 is -1 and \[(-1)^2=1\] so you can write \[x^2-2x=3\] and then \[(x-1)^2=3+1=4\] the last step was adding 1 to both sides so that the left hand side is now a perfect square

OpenStudy (anonymous):

then solve as \[(x-1)^2=4\] \[x-1=2\] or \[x-1=-2\] so \[x=3\] or \[x=-1\]